OpenStudy (anonymous):

Help ...Multiply (x^8+x^7)(x^8+x^7-x^6) would I just add the exponents in each? x^36???

OpenStudy (anonymous):

no because you have a sum

OpenStudy (anonymous):

you actually have to do 6 multiplications, and each of those you would add the exponents

OpenStudy (anonymous):

or is the answer this x^16+2x^15-x^13

OpenStudy (anonymous):

What? where is the sum?

OpenStudy (anonymous):

That is correct I think. Good job!

OpenStudy (anonymous):

\[(x^8+x^7)(x^8+x^7-x^6) \] \[x^8\times x^8+x^8\times x^7+x^8\times (-x^6)+x^7\times x^8+x^7\times x^7+x^7\times (-x^6)\]

OpenStudy (anonymous):

what you wrote is correct!

OpenStudy (anonymous):

Yay me! :) thanks guys!!!!

myininaya (myininaya):

\[(x^7)(x+1)(x^6)(x^2+x-1)=x^{7+6}(x+1)(x^2+x-1) \] \[=x^{13}(x^3+x^2-x+x^2+x-1)=x^{13}(x^3+2x^2-1)=x^{16}+2x^{15}-x^{13}\]

myininaya (myininaya):

gj!